OpenStudy (amonoconnor):

I am REALLY stuck on this one... This (I think) Definite Integral problem has me stuck at the first step, and I can usually at least get my "u" figured out... Would someone be willing to coach me through how to approach this problem, and solving it? *I will post the question as the first comment below. Any and all help is greatly appreciated!

OpenStudy (tkhunny):

Don't forget to show us YOUR efforts.

OpenStudy (amonoconnor):

The problem: \[\int\limits_{0}^{a}x \sqrt({x^2 + a^2})dx, a >0.\]

OpenStudy (mathmale):

Strongly suggest that you consider a substitution. What would that subst. be?

OpenStudy (amonoconnor):

I tried using (x^2 + a^2)^(1/2) as u, but got nowhere, and I can vaguely see some relationships in the book and WolframAlpha'a answers, but to be honest, the "a" is throwing me off, and I'm not sure where to start here... :/

OpenStudy (amonoconnor):

...Is the best "u" going to be x^2?

OpenStudy (tkhunny):

Maybe just the stuff inside the radical? All of it?

OpenStudy (mathmale):

No, but that's better than your previous venture.

OpenStudy (mathmale):

Let u = x^2 + a^2. Differentiate that, please. Find du/dx.

OpenStudy (mathmale):

Notice that u=x^2+a^2 is all INSIDE the parentheses. Do not include the exponent in your substitution u (at least, not in this problem).

OpenStudy (mathmale):

Hints: x usually represents a variable, whereas a usually represents a constant.

OpenStudy (amonoconnor):

Gotcha! So... \[u = x^2 + a^2\] \[du = (2x + a^2)dx\] ?

OpenStudy (tkhunny):

Maybe \(u^{2} = x^{2} + a^{2}\) Did you know you could do that?

OpenStudy (amonoconnor):

Nooo! :D So... u = x+1?

OpenStudy (mathmale):

Stop for a minute and think: why are we bothering with a substitution? What would be its purpose? Note that you have Radical (x^2+a^2), with x outside the radical.

OpenStudy (mathmale):

Before we go any further, consider that a and a^2 are not variables, but constants, using the conventions that x usually represents a variable and a a constant.

OpenStudy (tkhunny):

Say it with me, now..."The derivative of a constant is ______."

OpenStudy (amonoconnor):

To get a derivative that's in the integrand?

OpenStudy (mathmale):

what is the derivative of a constant? What is \[\frac{ d }{ dx }a^2?\]

OpenStudy (amonoconnor):

Oh, duh Amon.. they're 0.

OpenStudy (amonoconnor):

?

OpenStudy (mathmale):

Correct. So, if your u is x^2+a^2, du/dx is what?

OpenStudy (amonoconnor):

I thought you meant treat it like "e" or something.

OpenStudy (amonoconnor):

Then... \[du = 2x*dx\]

OpenStudy (mathmale):

e without an exponent is another constant. e with an exponent represents an exponential function.

OpenStudy (mathmale):

Very good. du=2xdx. Note that you do have x outside the radical, but not the 2. Is there nevertheless a way to simplify the integral using substitutions?

OpenStudy (amonoconnor):

Yes, I think? \[\frac{1}{2}du = x*dx\]

OpenStudy (tkhunny):

e - constant e^2 - constant e^x - exponential function

OpenStudy (amonoconnor):

Then we can plug in the left for the right, and make what's in the radical "u".

OpenStudy (mathmale):

Your original integral includes xdx, but not 2xdx. What are you gonna do about that? Good work ... you answered my qu as I was typing it. So, replace your xdx with du/2. Under the integral, replace your x^2+a^2 with u.

OpenStudy (mathmale):

What do you end up with? Is it any easier to integrate than the previous form? Let's hope so.

OpenStudy (mathmale):

Would appreciate your typing in the integral as it looks after 2 substitutions.

OpenStudy (amonoconnor):

So given the algebra we did earlier working with "u"... \[= \int\limits_{0}^{a}\frac{1}{2}\sqrt{u}*du\] \[= \frac{1}{2}[\frac{2}{3}u^{3/2}] 0/a\]

OpenStudy (amonoconnor):

\[= \frac{1}{3}(x^2 + a^2)^{3/2}, from 0 \to a\]

OpenStudy (mathmale):

The limits of integration AFTER substitution are not necessarily the same as BEFORE substitution. Note that y ou let u=x^2+a^2.

OpenStudy (amonoconnor):

But if I plug "u" back in, and am not plugging in the limit directly for the term "u" itself, then I keep the limits that were (are) in terms of "x" (or whatever the problem is initially in terms of), aye?

OpenStudy (mathmale):

You can do that, but after going thru the trouble of substituting, why not keep u and forget about x^2+a^2? It's so much simpler. If x goes from 0 to a, what happens to u? u is your new variable and you thus need to compute the appropriate limits on u.

OpenStudy (mathmale):

There's more than one way to skin a cat. You can either STICK with u, or you could re-substitute x^2+a^2 for u. My preference is that you stick with u.

OpenStudy (mathmale):

x varies from 0 to a. Thus, if u=x^2+a^2, u varies from (what?) to (what?)

OpenStudy (mathmale):

The latter values are your new limits of integration./

OpenStudy (amonoconnor):

Gotcha:) I will agree to respectfully disagree with you, as you are far more knowledgable than I, but I will plug x^2 + a^2 back in:) Umm, u varies from... [2a^2 , a^2] ???

OpenStudy (mathmale):

You can do that substitution; it's correct. However, by going back to x^2+a^2, you have abandoned u.

OpenStudy (mathmale):

Do it both ways. You choose to plug x^2+a^2 back into the RESULT of your integration? Fine. Then let x vary from 0 to a. You cvhoose to stick with u? Fine. Since u=x^2+a^2, and a varies from 0 to a, u varies from a^2 to 2a^2.

OpenStudy (mathmale):

I admire hunny's suggestion, that you let u^2=x^2+a^2. Doing this will eliminate the square root operation, which is a good thing. Simplification!

OpenStudy (mathmale):

Regardless of the method used, you MUST end up with the same result. Else something is wrong.

OpenStudy (amonoconnor):

I'm with you there! I'm working it out now... :)

OpenStudy (amonoconnor):

Okay, I got: \[ = \frac{ 1 }{ 3 }(a^2)^{3/2}\] ??

OpenStudy (amonoconnor):

Is that right?

OpenStudy (mathmale):

Question: Can you simplify (a^2)^(3/2)? making good progress. I have to get off the 'Net right now, but will be back in 0.5-1.0 hour. Good luck!

OpenStudy (amonoconnor):

Yes, you can, but is that the right answer in general, if I make (a^2)^(3/2) = a^(7/2)?

OpenStudy (mathmale):

Let's re-write (a^2)^(3/2) as\[(a^2)^{\frac{ 3 }{ 2 }}\]

OpenStudy (mathmale):

since one of the rules of exponentiaion is (x^a)^b=x^(ab), \[(a^2)^{\frac{ 3 }{ 2}}=\]

OpenStudy (mathmale):

what?

OpenStudy (amonoconnor):

Where is WolframAlpha getting the sqrt. of 2 here?

OpenStudy (amonoconnor):

My bad on the exponents there...

OpenStudy (mathmale):

If you use substitution, the integration becomes quite simple. I'd suggest you go thru with the integration yourself and then worry about that Sqrt(2) from Wolfram.

OpenStudy (amonoconnor):

Here's what I got: \[\frac{1}{3}[x^2 + a^2]^{3/2}, 0/a\] \[= [\frac{1}{3}((a)^2+a^2)^{3/2}] - [\frac{1}{3}((0)^2+a^2)^{3/2}]\] \[= [\frac{1}{3}(2a^2)^{3/2}] - [\frac{1}{3}(a^2)^{3/2}]\] \[= [\frac{1}{3}(a^2)^{3/2}] = \frac{ 1 }{ 3 }a^{3}\] How does that look? :/

OpenStudy (mathmale):

Sorry: I was detracted by another posted problem.

OpenStudy (mathmale):

Having taken a quick (not really careful) look at your work, my first impression is that you've done well.

OpenStudy (mathmale):

Cool. Have to get off the 'Net pronto, but will be back later. Admire your persistence and hard work! See you later.

OpenStudy (amonoconnor):

Do you guys know where the \[2\sqrt{2} - 1 \] comes from? :/

OpenStudy (freckles):

\[= [\frac{1}{3}(2a^2)^{3/2}] - [\frac{1}{3}(a^2)^{3/2}]\] so you went from here to: \[= [\frac{1}{3}(a^2)^{3/2}] = \frac{ 1 }{ 3 }a^{3}\] how did you make the 2^(3/2) disappear?

OpenStudy (freckles):

\[=\frac{1}{3}2^\frac{3}{2}a^3-\frac{1}{3}a^3 \\ =\frac{1}{3}a^3(2^\frac{3}{2}-1)\]

OpenStudy (amonoconnor):

I thought that the (2a^2)^(3/2) - (a^2)^(3/2) was just like subtracting 5x from 6x. So basically, the issue there was that I needed to distribute the power of 3/2 to the 2, and not treat it as such a coefficient?

OpenStudy (freckles):

you could have treated like if was outside the exponent if it was outside the exponent :p

OpenStudy (freckles):

\[2(a^2)^\frac{3}{2}-(a^2)^\frac{3}{2}=(2-1)(a^2)^\frac{3}{2}=1(a^2)^\frac{3}{2} \\ \text{ but } (2a^2)^\frac{3}{2}-(a^2)^\frac{3}{2}= 2^\frac{3}{2}(a^2)^\frac{3}{2}-(a^2)^\frac{3}{2}=(a^2)^\frac{3}{2}(2^\frac{3}{2}-1)\]

OpenStudy (amonoconnor):

Ah.... I see.. :) Gotcha.

OpenStudy (freckles):

example (2*5)^2 is not the same as 2*(5^2)

OpenStudy (freckles):

(2*5)^2=10^2=100 or you could say (2*5)^2=2^2*5^2=4*25=100 but 2*(5^2)=2*25=50

OpenStudy (freckles):

2*(5^2) or 2*(5)^2 whatever

OpenStudy (freckles):

or you can even say 2*5^2 :p

OpenStudy (freckles):

2*(5^2)=2*(5)^2=2*5^2 but none of those are equal to (2*5)^2

OpenStudy (amonoconnor):

So how would you treat or simplify this then: \[[\frac{1}{3}(a^2 + a^2)^{3/2}] = [\frac{1}{3}(2a^2)^{3/2}] ????\]

OpenStudy (amonoconnor):

That's where I pulled the two into play, but if it's not to be treated like a coefficient, I'm confused about what's going on. :?

OpenStudy (freckles):

the 3/2 is on the (2a^2)

OpenStudy (freckles):

\[\frac{1}{3}(2a^2)^\frac{3}{2} \\ \frac{1}{3}(2^1 a^2)^\frac{3}{2} \\ \frac{1}{3}(2^{1 \cdot \frac{3}{2}} a^{2 \cdot \frac{3}{2}})\]

OpenStudy (amonoconnor):

right, so doesn't it distribute into the "^2", just as it would if it was for example (x^2)^(3/2)?

OpenStudy (amonoconnor):

Okay, I see where you're going there, but that's just so not intuitive for me... My bad I guess!

OpenStudy (freckles):

\[a^{2 \cdot \frac{3}{2}}=a^3\] since a>0

OpenStudy (amonoconnor):

Right, I just didn't get the whole distributing the power across to the 2 in front of that a^2. :) I see what you mean completely now, albeit being still in a phase of working adjusting to thinking of everything this way. Thank you:)

OpenStudy (freckles):

oh ok well you could just ignore 2 being raised to the 3/2 power

OpenStudy (freckles):

\[\text{ for example } 2^3=8 \text{ \not } 2 \]

OpenStudy (freckles):

I know that seems like a stupid example \[(2a)^n=2^n a^n\]

OpenStudy (amonoconnor):

Gotcha... gotcha indeed. I learn something new every night... :)

OpenStudy (freckles):

lol and you probably will for the rest of your life we never stop learning well until we become a vegetable or dead

OpenStudy (amonoconnor):

Good Lord, I better never stop learning!!!!

OpenStudy (freckles):

lol

OpenStudy (freckles):

good luck with living forever

OpenStudy (amonoconnor):

;) I'll be here!